# Quantifying the Dunkelflaute: An analysis of variable renewable energy droughts in Europe

### v2

@@ 2.2 Frequency-duration distribution: short drought events occur much more often than longer ones
A cumulative frequency-duration distribution shows how often VRE droughts that lasted at least a certain duration occurred on average per year in the investigated data. Figure 2 illustrates these distributions across all investigated relative thresholds for Germany and Spain, as examples of central and southern European countries with different wind and solar energy resources, and for the European copperplate scenario. For a given threshold, the frequencies of all events that are at least as long as a given duration are shown in ascending order on the vertical axis. For instance, at a relative threshold , Germany experienced on average 9.8 onshore wind droughts lasting at least 48 hours (two days), 5.2 lasting at least 168 hours (one week), and 3.1 lasting at least 336 hours (two weeks) per year. Note that the duration axis in Figure 2 is truncated at 360 hours, so it focuses on the occurrence of shorter-duration events. These are particularly relevant for the operation of shorter-duration flexibility options in the power sector, e.g., battery or hydro reservoir storage. Longer-lasting drought exceeding 360 hours, which may drive the need for long-duration electricity storage or other firm low-carbon technologies, are not visible. The following sections focus more on such extreme periods.
@@
The Figure also illustrates the geographical balancing effect, as droughts occur less often in an ideal European interconnection than in single countries. This is true for both single VRE technologies and portfolios. The reason for this is that periods of low solar and wind availability are not perfectly correlated across Europe [roth_geographical_2023]. For example, assuming a relative threshold , there have been on average approximately eight and five VRE portfolio droughts per year in Germany and Spain that lasted at least one week, respectively, while less than four of such events occurred in Europe.
In the SI, we provide additional results on the seasonality of drought patterns for Europe under the assumption of perfect interconnection. For low thresholds and events lasting longer than a few days, PV droughts are more frequent in winter, while wind droughts are more frequent in summer (Figure SI.7). The SI also shows return periods, which are the reciprocals of yearly frequencies. Similar to frequencies, return periods vary substantially between technologies and regions (see Section SI.4.)
@@ 2.3 Maximum drought duration strongly depends on threshold and differs across years and seasons
For medium thresholds, we find that both a technology portfolio and geographical balancing strongly decrease the maximum drought duration. For example, for a relative threshold of , the longest PV drought in Germany lasts 291 days, while the maximum on- and offshore wind droughts are 303 and 209 days, respectively. In the case of a combined VRE portfolio, the respective maximum drought in Germany decreases to 106 days due to the technology portfolio effect. If, on top, perfect interconnection in Europe is considered, the duration of the longest portfolio event for the same threshold further declines to 55 days due to the geographical balancing effect.
@@
Especially for medium thresholds, the maximum yearly drought duration strongly depends on the weather year (Figure SI.9). For example, the longest VRE portfolio drought in Germany obtained for a relative threshold of occurred in 1996 (106 days), while the shortest yearly maximum drought duration was in 2018 (22 days). Considering perfect interconnection across all countries and the same threshold, the longest and shortest portfolio droughts occurred in 1997 and 1999, respectively (55 and 8 days).
@@
The most extreme droughts vary substantially across countries, both with respect to their duration and the year of occurrence (Figures SI.10 and SI.11). Importantly, the ranking of years varies with the drought threshold (Figure 3). For instance, the longest VRE portfolio drought in Germany for a threshold of occurred in 2007 but for the thresholds and the most extreme droughts were in 1997 and 2003. The corresponding longest drought events in Spain (or in Europe, assuming perfect interconnection) occurred in 1991 (2005), 2017 (1996), and 2017 (2002). The spreads between the longest drought events in each year further increase with higher thresholds, before they decrease again approaching the threshold (Figure SI.12). From these findings follows that any claim on a specific year being particularly relevant for renewable drought analyses should always be qualified with the respective threshold.
@@
The maximum drought duration further varies substantially across seasons and technologies. Solar PV droughts primarily occur in fall and winter, with their median hour in November, December, or January (Figure 4). The maximum European PV drought under unconstrained geographical balancing lasts 248 days with a median hour in December for a threshold of . The longest PV droughts with median hours in spring or summer months barely last longer than a few days, except for those identified by very high thresholds close or equal to . In contrast, the longest European on- and offshore wind droughts mainly occur in spring and summer, with medians in July and maximum duration of 218 or 204 days, respectively. Maximum wind drought duration in winter months are much lower. In single countries, summer and winter wind droughts can be significantly longer (balancing effect). The seasonal complementarity of wind and solar further mitigates combined technology portfolio droughts to a substantial extent (portfolio effect). The longest-lasting droughts occur during winter months, which is a result of low seasonal PV availability in combination with rare but severe concurrent wind droughts. The median hours of the longest renewable technology portfolio droughts in Germany (106 days) or Spain (137 days) occur in November, or, under the assumption of unconstrained balancing, in January (55 days).
@@ 2.4 Portfolio and balancing effects mitigate maximum drought durations
Comparing weighted averages over all thresholds and countries, the maximum drought duration of a VRE portfolio decreases by 64%, 52%, or 47% compared to solar PV, onshore wind, or offshore wind alone (Table 1). This portfolio effect is more pronounced for PV than for wind power, as onshore and offshore wind power complement both diurnal and seasonal shortages of PV. In contrast, wind power does not incur diurnal shortages. Assuming a European copperplate (CP), the portfolio effect (-80%) for solar PV is even stronger than for any isolated country. This is because in a larger geographical area, there is more complementary wind power available to compensate for solar PV shortages [roth_geographical_2023]. In turn, the portfolio effect is less pronounced for onshore wind power in the European interconnection (-45%) as PV availability is more homogeneously distributed. For offshore wind power, the portfolio effect is higher (-70%), as only a subset of countries features this technology, such that solar and wind onshore resources of all countries can compensate for extreme offshore droughts in a subset of countries.

### v3

@@ 2.2 Frequency-duration distribution: short drought events occur much more often than longer ones
A cumulative frequency-duration distribution shows how often VRE droughts that lasted at least a certain duration occurred on average per year in the investigated data. Figure 2 illustrates these distributions across all investigated relative thresholds for selected regions. For a given threshold, the frequencies of all events that are at least as long as a given duration are shown in ascending order on the vertical axis. For instance, at a relative threshold , Germany experienced on average 9.8 onshore wind droughts lasting at least 48 hours (two days), 5.2 lasting at least 168 hours (one week), and 3.1 lasting at least 336 hours (two weeks) per year. For comparability, Figure 2 uses the same frequency-axis range across all technologies and regions. Figure SI.7 focuses on on- and offshore wind droughts using an adjusted frequency-axis range. Also note that the duration axis in Figure 2 is truncated at 360 hours, so it focuses on the occurrence of shorter-duration events. These are particularly relevant for the operation of shorter-duration flexibility options in the power sector, e.g., battery or hydro reservoir storage. Figure SI.8 shows longer-lasting drought exceeding 360 hours to up to one full year, which may drive the need for long-duration electricity storage or other firm low-carbon technologies. Additionally, the following sections analyze such extreme periods in depth.
Cumulative frequency-duration distributions of drought events across all investigated relative drought thresholds with , sorting the yearly frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific yearly frequency. For illustration, the distributions are truncated at 360 hours, i.e., they show events with a maximum duration of just above two weeks. Yearly frequencies of events lasting at least two days, one week, and a fortnight are marked for a relative threshold of . An interactive version of this Figure which allows zooming and rotating is available online [kittel_high_2025].
@@
Figure 2 also illustrates the geographical balancing effect, as droughts occur less often in an ideal European interconnection than in single countries. This is true for both single VRE technologies and portfolios. The reason for this is that periods of low solar and wind availability are not perfectly correlated across Europe [roth_geographical_2023]. For example, assuming a relative threshold , there have been on average approximately eight and five VRE portfolio droughts per year in Germany and Spain that lasted at least one week, respectively, while less than four of such events occurred in Europe. For on- and offshore wind, the balancing effect mitigates drought frequency, particularly for drought events identified by lower thresholds, i.e., more severe wind droughts. This effect is visualized by the additional white space in Figure 2 and the frequency-duration distributions in Figure SI.7 when in the European copperplate scenario compared to Germany or Spain, respectively.
In the SI, we provide additional results on the seasonality of drought patterns for Europe under the assumption of perfect interconnection. For low thresholds and events lasting longer than a few days, PV droughts are more frequent in winter, while wind droughts are more frequent in summer (Figure SI.9).
2.3 Return periods: most extreme events occur rarely
Return periods of VRE droughts are given by the reciprocal of average yearly frequencies. Return-period curves for extreme events indicate the longest event duration that can, on average, be expected to reoccur after a given number of years (Figure 3). Note that the return period monotonically increases and the maximum return period is limited by the 38 years of data that we investigate. In SI.4, we provide more detailed, three-dimensional return-period visualizations.
For each threshold, the maximum event duration increases with higher return periods, but is highly sensitive to the underlying threshold. For any given return period, higher thresholds lead to events with longer maximum durations, plateauing at 365 days for very high thresholds (Figure SI.10). Note that such high durations indicate below-average wind and solar years and should not be mistaken for actual drought events. The return period-duration curves vary substantially between technologies and countries. For example, the maximum offshore wind drought for a relative threshold of in Germany that can be expected to return every 20 years lasts 198 days and is much shorter than the corresponding PV (289 days) and onshore wind droughts (297 days). By comparison, in Spain, such 20-year return period droughts are shorter for PV (216 days) and onshore wind (245 days) but longer for offshore wind (221 days).
When combined in a technology portfolio, the maximum duration decreases substantially for each return period in both Germany and Spain and also in the European copperplate, which is another instance of the portfolio effect. The balancing effect is also visible: with perfect interconnection, the maximum drought duration further decreases due to a limited temporal correlation of wind droughts across Europe, especially for the renewable technology portfolio. In the case of solar PV, higher solar availability in South Europe during winter balances lower availability in Northern Europe.
Return period curves of extreme drought events with an average annual frequency below 1 across selected relative thresholds . Figure SI.10 provides a more comprehensive perspective on return periods across all thresholds.
@@ 2.4 Maximum drought duration strongly depends on threshold and differs across years and seasons
Most extreme duration of single drought events across all investigated years for given relative thresholds with . The year with the most extreme event duration varies across thresholds. The marked numbers indicate the longest drought durations found for a relative threshold of .
@@
For medium thresholds, we find that both a technology portfolio and geographical balancing strongly decrease the maximum drought duration. For example, for a relative threshold of , the longest PV drought in Germany lasts 291 days, while the maximum on- and offshore wind droughts are 303 and 209 days, respectively. In the case of a combined VRE portfolio, the respective maximum drought in Germany decreases to 106 days due to the technology portfolio effect. If, on top, perfect interconnection in Europe is considered, the duration of the longest portfolio event for the same thresh old further declines to 55 days due to the geographical balancing effect. As visible in Figure 4, the balancing effect is particularly pronounced for medium thresholds. This is because maximum drought durations are very short for very small thresholds close to 0 even without interconnection, and very long for large thresholds approaching 1 even with perfect interconnection. Accordingly, the balancing effect cannot play out for such extreme thresholds.
@@
Especially for medium thresholds, the maximum yearly drought duration strongly depends on the weather year (Figure SI.11). For example, the longest VRE portfolio drought in Germany obtained for a relative threshold of occurred in 1996 (106 days), while the shortest yearly maximum drought duration was in 2018 (22 days). Considering perfect interconnection across all countries and the same threshold, the longest and shortest portfolio droughts occurred in 1997 and 1999, respectively (55 and 8 days).
@@
The most extreme droughts vary substantially across countries, both with respect to their duration and the year of occurrence (Figures SI.12 and SI.13). Importantly, the ranking of years varies with the drought threshold (Figure 4). For instance, the longest VRE portfolio drought in Germany for a threshold of occurred in 2007 but for the thresholds and the most extreme droughts were in 1997 and 2003. The corresponding longest drought events in Spain (or in Europe, assuming perfect interconnection) occurred in 1991 (2005), 2017 (1996), and 2017 (2002). The spreads between the longest drought events in each year further increase with higher thresholds, before they decrease again approaching the threshold (Figure SI.14). From these findings follows that any claim on a specific year being particularly relevant for energy modeling analyses, because it features a large renewable drought, should always be qualified with the respective threshold.
Most extreme duration of single drought events across all months in the data and all investigated thresholds with . The contour lines represent the threshold-month-specific maximum duration. Note that the monthly assignment illustrates the median hour of respective droughts, while the duration of each event is plotted on the vertical axis. Events lasting longer than one month start (end) in previous (subsequent) months. The events with the highest and lowest duration across all months are marked for a threshold .
@@
The maximum drought duration further varies substantially across seasons and technologies. Solar PV droughts primarily occur in fall and winter, with their median hour in November, December, or January (Figure 5). The median hour represents the hour after half of the drought duration has passed. The maximum European PV drought under unconstrained geographical balancing lasts 248 days with a median hour in December for a threshold of . The longest PV droughts with median hours in spring or summer months barely last longer than a few days, except for those identified by very high thresholds close or equal to . In contrast, the longest European on- and offshore wind droughts mainly occur in spring and summer, with medians in July and maximum duration of 218 or 204 days, respectively. Maximum wind drought duration in winter months are much lower. In single countries, summer and winter wind droughts can be much longer (balancing effect). The seasonal complementarity of wind and solar further mitigates combined technology portfolio droughts to a substantial extent (portfolio effect). The longest-lasting droughts occur during winter months, which is a result of low seasonal PV availability in combination with rare but severe concurrent wind droughts. The median hours of the longest renewable technology portfolio droughts in Germany (106 days) or Spain (137 days) occur in November, or, under the assumption of unconstrained balancing, in January (55 days).
@@ 2.5 Portfolio and balancing effects mitigate maximum drought durations
Reduction of maximum drought durations for single renewable technologies compared to a renewable technology portfolio within a country (portfolio effect) and when considering unconstrained geographical balancing across all countries (balancing effect) averaged over all thresholds and countries.
@@
Comparing weighted averages over all thresholds and countries, the maximum drought duration of a VRE portfolio decreases by 64%, 52%, or 47% compared to solar PV, onshore wind, or offshore wind alone (Figure 6). This portfolio effect is more pronounced for PV than for wind power, as onshore and offshore wind power complement both diurnal and seasonal shortages of PV. In contrast, wind power does not incur diurnal shortages. Assuming a European copperplate (CP), the portfolio effect (-80%) for solar PV is even stronger than for any isolated country (Table SI.1). This is because in a larger geographical area, there is more complementary wind power available to compensate for solar PV shortages [roth_geographical_2023]. In turn, the portfolio effect is less pronounced for onshore wind power in the European interconnection (-45%) as PV availability is more homogeneously distributed. For offshore wind power, the portfolio effect is higher (-70%), as only a subset of countries features this technology, such that solar and wind onshore resources of all countries can compensate for extreme offshore droughts in a subset of countries.